A given line has a slope of 3/4 and passes through a point (-2, 5). Find the equation of the line?
and
Given f(x) = |3x - 2|, find the value of f(-4) + f(-1).
(answer is either 15 or 19)
I/m stuck on these, think I know but not sure.

Answer by rfer(16322)

A)
5=3/4*-2+b
5=-6/4+b
5=-1.5+b
b=6.5
-
y=3/4x+6.5